# The Low-Temperature Phenomenology of Gap Inhomogeneity in the Cuprate Superconductors: High-Energy Granularity, Low-Energy Homogeneity, and Spectral Kinks

### v1

@@ IV Intermediate-Sized Patches ($l\approx\xi$)
We had noted earlier that the correlation between $\Delta_{T}$ and $\Delta_{S}$ as evidenced from the real-space plots in Fig. 8 is not perfect. More pictorial evidence for this comes from comparing these two quantities along the same linecut (Fig. 11b). $\Delta_{S}$ generally tracks $\Delta_{T}$, but we can observe that for sites where $\Delta_{T}$ is very small, the corresponding $\Delta_{S}$ is always larger and offset to the right. In contrast, for large-$\Delta_{T}$ sites, these two quantities match each other well. That said, $\Delta_{S}$ exhibits slower spatial fluctuations than $\Delta_{T}$ along the same linecut; rapid variations in $\Delta_{T}$ are not reflected by a similar amount of variation in $\Delta_{S}$, and these two quantities match each other best at sites lying within clusters of multiple patches with similar values of $\Delta_{T}$ are present. As for $\Delta_{L}$, we find that not only is its variation much smaller than either $\Delta_{T}$ or $\Delta_{S}$, but its variations are also much more long-ranged than the other two quantities: while $\Delta_{T}$ and $\Delta_{S}$ vary on the scale set by $l$, $\Delta_{L}$ on the other hand varies over longer length scales (several multiples of $l$). The sharp boundaries are largely invisible to $\Delta_{L}$, which only curves slightly in the presence of a boundary between small- and large-$\Delta_{S}$ regions. We also find that $\Delta_{L}$ tends to diverge from $\Delta_{T}$ or $\Delta_{S}$ most strongly when $\Delta_{T}$ is large; in this instance, it tends to remain near $\Delta_{L}\approx\overline{\Delta}$, and only increases mildly in the presence of this strong-pairing patch.
@@
In sum, what we have observed here is that for the inhomogeneous $d$-wave superconductor with $l\approx\xi$, the low-energy ($\omega<\Delta_{K}$) spectra are homogeneous (for the $l=8$ results here, the low-energy spectra are that of a homogeneous $d$-wave superconductor with $\Delta=0.16$), while the high-energy ($\omega>\Delta_{K}$) LDOS exhibits much of the full inhomogeneity of the underlying order parameter, but with uneven corresponences between $\Delta_{S}$ and $\Delta_{T}$. Notably, proximity effects are strong; this is most evident in the spectra with $\Delta_{T}\approx 0$, which show a finite $\Delta_{S}$. The critical energy $\Delta_{K}$, which separates homogeneous and inhomogeneous states, in turn appears to be largely set by the smallest spectral gaps in the system. A crucial difference between the $l=8$ regime studied here and the $l=1$ one in the previous section is that the quasiparticle response of the system to order-parameter inhomogeneity largely hews to the underlying order parameter at high energies such that $\Delta_{T}$ and $\Delta_{S}$ are positively correlated, but at low energies the response reflects little of the order-parameter variation (so that $\Delta_{L}$ is only weakly correlated with $\Delta_{T}$). Evidently, increasing the inhomogeneity length scale $l$ reduces the energy range in which the spectrum is homogeneous, so that for sufficiently large $l$, the resulting critical energy scale $\Delta_{K}$ separating homogeneous and inhomogeneous spectra is pushed downwards towards the Fermi energy.
@@ V Large Patches ($l>\xi$)
Figure 14: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (b), and $\Delta_{L}$ (c), and the absolute value of their respective Fourier transforms (d, e, and f) for patch size $l=32$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
Figure 15: Left: Plots of the binned average local density of states for $l=32$. Each LDOS spectrum is binned according to its spectral gap (with 10 bins in all), and all spectra in a given bin are then averaged over. Middle: Plot of the distribution of the local density of states as a function of energy (heat map) and the average local density of states (dashed white line) for $l=32$. The spectra are binned according to the energy, and histograms for each energy are taken. The counts per energy bin are shown as a heat map. For both these plots, 4 disorder realizations and a total of 1,048,576 individual spectra are used in this plot. Right: Plots of the density of states for a clean bulk $d$-wave superconductor for various values of $\Delta$. To obtain these plots, we made use of a $1024\times 1024$ momentum-space grid, and the same band-structure and broadening parameters in the inhomogeneous cases are used here.
@@
In this regime, both $\Delta_{T}$ and $\Delta_{S}$ maps look mostly identical to one another; this can be seen in Fig. 14a and b. The difference is primarily in the more rounded shape of the patches in $\Delta_{S}$, and in the fact that when adjacent patches in $\Delta_{T}$ have similar (but not identical) values, they show up in $\Delta_{S}$ as a large, mostly uniform patch, with the gradations in $\Delta_{T}$ not visible in $\Delta_{S}$. Also, isolated small-$\Delta_{S}$ regions tend to be slightly larger in spatial extent than similarly isolated large-$\Delta_{S}$ regions, an observation that is also clearly seen in the $l\approx\xi$ case; this is a reflection of the larger local coherence length within these small-gap patches, which ensures that (through the proximity effect) the effect of the small-$\Delta_{T}$ patch “leaks out” into the surrounding large-$\Delta_{T}$ region. Conversely, a large-$\Delta_{T}$ patch surrounded by a small-$\Delta_{T}$ would feature the “leakage” of the small-gap background into the patch, making the large-$\Delta_{S}$ patch look smaller than its actual size in $\Delta_{T}$ is. This effect for $l>\xi$ however is much more muted compared to what is seen in the $l\approx\xi$ case. Finally, similar to the $l=1$ and $l=8$ cases discussed earlier, the $\Delta_{S}$ map always is nonzero, with a minimum $\Delta_{S}$ present even in regions where $\Delta_{T}$ is close to zero. Even with $l=32$, proximity effects are strong enough that a nonzero pair density, and consequently a finite gap, is present deep within these patches. This minimum gap however is smaller than that seen in the $l\approx\xi$ case.
@@
The map of the low-energy gap $\Delta_{L}$ with $l=32$ (Fig. 14c) now clearly reflects the inhomogeneous nature of the underlying order parameter. Small- and large-$\Delta_{T}$ regions match up neatly with small- and large-$\Delta_{L}$ regions, and this time the distribution of $\Delta_{T}$ is broad enough to capture the variations in $\Delta_{T}$ to reasonable accuracy. The agreement is clearly not perfect, though: the finer gradations seen in $\Delta_{T}$ are not captured by $\Delta_{L}$, and the distribution of values of $\Delta_{L}$ still appears to be narrower than that of $\Delta_{S}$. Nevertheless, the $\Delta_{L}$ plots demonstrate that for $l>\xi$, the low-energy states now display more inhomogeneity than those in the $l\approx\xi$ and $l<\xi$ cases, but still not as much as the high-energy states do. The overall similarity among the $\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$ maps carries over to their respective Fourier transforms (Figs. 14d, e and f). In particular, the Fourier transforms of the $\Delta_{T}$ and $\Delta_{S}$ maps look almost identical, with both featuring a highly localized region of high intensity near zero wavevector and higher harmonics arising from the regular square shape of the patches in both real-space maps. The Fourier transform of the $\Delta_{L}$ map also features strong intensities only at small wavevectors. One can see that the small-momentum part of Fig. 14f is similar to that of Figs. 2f and 8f, but as higher-momenta scattering processes are suppressed for the $l=32$ case, the only feature in the Fourier transform left is a much smaller X-shaped feature near zero momentum.
@@
Similar to the $l=8$ case, at $\omega=0.16$ (Fig. 16c), there are three kinds of regions: large-LDOS patches which correspond to patches where the local spectral gap $\Delta_{S}\approx 0.16$; small-LDOS patches which correspond to large-$\Delta_{S}$ regions with $\Delta_{S}>0.16$; and regions hosting prominent modulations which correspond to small-$\Delta_{S}$ regions with $\Delta_{S}<0.16$. These striking modulations are (like the ones seen in the $l=8$ case) a result of quasiparticle scattering off of the walls of the order-parameter barrier formed by the large-$\Delta_{T}$ neighboring patches. These regions with modulations grow in size with increasing $\omega$. At $\omega=0.31$ (Fig. 16d), the only regions without any visible modulations are the large-LDOS patches corresponding to patches with large $\Delta_{S}$; these modulations are present everywhere else. A peculiarity of our square-patch disorder model is that when $l$ is large (such as in the present $l=32$ case), the boundaries of the small-$\Delta_{S}$ regions that host these modulations are straight and have right-angle corners, which result in mostly ordered modulations in the $(1,0)$ and $(1,0)$ directions. This is in contrast to that seen in the $l=8$ case earlier where, as a result of the small value of $l$, the regions that have these modulations are irregularly shaped and therefore the resulting scattering patterns are not as sharply defined as in the present case.
@@
To succinctly summarize all these observations together, we show once more the various two-dimensional histograms of the order-parameter measures in Fig. 18. The corresponding plot for the $\Delta_{T}$-$\Delta_{S}$ pair (Fig. 18a) shows that the two quantities are now very strongly correlated with one another, with only minimal fuzziness in the correspondences. It is worth noting that the system still exhibits a finite $\Delta_{S}$ everywhere, even in regions where $\Delta_{T}\approx 0$. Here, the smallest $\Delta_{S}\approx 0.05$, which is smaller than the smallest $\Delta_{S}$ for the $l\approx\xi$ case. This is consistent with the observation made earlier that the kink energy $\Delta_{K}$, which is the scale that separates homogeneous and inhomogeneous states and which looks to be tied to the smallest spectral gaps in the system, is pushed down in energy as the inhomogeneity length scale $l$ is increased. The histogram for the $\Delta_{T}$-$\Delta_{L}$ pair (Fig. 18b) now shows a much more pronounced positive correlation between these two quantities compared to the two cases considered before, with the agreement best at small $\Delta_{T}$ and worst at large $\Delta_{T}$. Finally, the $\Delta_{S}$-$\Delta_{L}$ pair (Fig. 18c) also appears to be more correlated with one another than in the previous cases. As with the $l\approx\xi$ case, here we find a similar tendency for the joint distribution to be mostly localized within one small region for small $\Delta_{S}$ and for it to extend out for increasing $\Delta_{S}$.
@@ VI Low-Energy Homogeneity and Spectral Kinks
Figure 20: Plots of the binned average local density of states for $l=1$ (left), $l=8$ (middle), and $l=32$ (right), with finer bins used than in the previous plots. Each LDOS spectrum is binned according to its spectral gap (with 41 bins in all), and all spectra in a given bin are then averaged over. For these plots, 4 disorder realizations and a total of 1,048,576 individual spectra are used. The width of each line plot is proportional to the number of spectra within the bin. Only lineshapes corresponding to bins with total number of spectra greater than 1024 are shown here.
@@
Perhaps the most notable behavior seen in STS experiments on the cuprates is the sharp dichotomy between the low-energy homogeneous states and the high-energy inhomogeneous states that are characteristic of the $l\gtrapprox\xi$ case, which is accompanied by the presence of kinks appearing in large-$\Delta_{S}$ spectra. We had noticed in Figs. 9a and 15a that these kinks tend to appear at a characteristic kink energy $\Delta_{K}$ which is roughly set by the smallest spectral gaps of the system. In this section, we will define $\Delta_{K}$ appropriately and study it as a function of $l$, contrast its behavior with the other energy scales that can be obtained from the LDOS spectra ($\Delta_{S}$ and $\Delta_{L}$), and put on quantitative footing the informal observations made in the previous section.
We first define $\Delta_{K}$ in quantitative terms. One can observe in Figs. 11a and 17a that a purely local definition of $\Delta_{K}$ as the energy where a kink is present is difficult to implement in practice in our numerics: at low energies, numerous wiggles in the LDOS as a function of energy can be seen, and it is difficult to ascertain if these are the true kinks that separate the homogeneous states from the inhomogeneous ones, or if these arise from finite-size effects. This difficulty of isolating the kink in the spectrum is otherwise not present in our operational definitions of $\Delta_{S}$ (which involves simply finding the peak of the positive-energy LDOS) and $\Delta_{L}$ (which involves taking the slope of the LDOS near $\omega=0$).
To get around this difficulty, we work instead with the bin-averaged spectra, as visualized in Fig 20. Here, we bin the individual spectra by its local spectral gap. Unlike Figs. 3a,  9a, and 15a, however, here we use much smaller bin widths that allow us to capture the averaged spectrum for each value of $\omega$. Because of the very large number of individual spectra, even a finer binning procedure allows for enough spectra per bin to ensure that an average over a sufficiently large number of spectra is performed for each individual bin. The averaging over a large number of spectra allows us to smoothen out out any finite-size effects, and we can thus define $\Delta_{K}$, as well as $\Delta_{S}$ and $\Delta_{L}$, for each of the bin-averaged spectra. $\Delta_{K}$ is defined as the lowest energy at which the derivative with respect to $\omega$ of the positive-energy bin-averaged LDOS exhibits a peak. Note that this definition implies that $\Delta_{K}\approx\Delta_{S}$ in the absence of a kink. Right away, we can expect the following behaviors to hold at these two extreme limits: a) in the extreme pointlike limit $l\ll\xi$ where the states are highly homogeneous, all three energy scales coincide, i.e., $\Delta_{S}\approx\Delta_{L}\approx\Delta_{K}\approx\text{constant}$; and b) in the extreme bulk limit $l\to\infty$, $\Delta_{S}\approx\Delta_{T}$ while $\Delta_{K}\to 0$.
@@
These results suggest that in the physically relevant $l\approx\xi$ regime, the notion of an average spectral gap—defined either by taking both the system- and configuration-average of $\Delta_{S}$ or by binning the LDOS and obtaining the bin-averaged $\Delta_{S}$, as we have done here—leads to a considerable overestimate of the true average underlying order parameter $\overline{\Delta}$. This is partly due to the finite $\Delta_{S}$ even in regions where $\Delta_{T}\approx 0$: these finite $\Delta_{S}$ values push up the average, leading to an overestimate of $\overline{\Delta}$. As we had noted earlier, these particular $\Delta_{S}$ values (i.e., the smallest gaps) are correlated with the value of $\Delta_{K}$; therefore, as $l$ increases, $\Delta_{K}$, and therefore these particular $\Delta_{S}$’s, decrease towards zero, and the average of $\Delta_{S}$ regresses to $\overline{\Delta}$ for sufficiently large $l$. It is telling that in Fig. 22, even patches as large as $l=64$ for the $\overline{\Delta}=0.32$ case still feature a large enough discrepancy between the averages of $\Delta_{S}$ and $\Delta_{L}$ to indicate that the system in this case is still far from the extreme bulk limit. On the other hand, taking the average of $\Delta_{L}$ appears to give a more accurate estimate of $\overline{\Delta}$, owing to the robustness of low-energy homogeneity when $l\approx\xi$.

### v2

@@ IV Intermediate-Sized Patches ($l\approx\xi$)
We had noted earlier that the correlation between $\Delta_{T}$ and $\Delta_{S}$ as evinced from the real-space plots in Fig. 7 is not perfect. More pictorial evidence for this comes from comparing these two quantities along the same linecut (Fig. 10b). $\Delta_{S}$ generally tracks $\Delta_{T}$, but we can observe that for sites where $\Delta_{T}$ is very small, the corresponding $\Delta_{S}$ is always larger and offset to the right. In contrast, for large-$\Delta_{T}$ sites, these two quantities match each other well. That said, $\Delta_{S}$ exhibits slower spatial fluctuations than $\Delta_{T}$ along the same linecut; rapid variations in $\Delta_{T}$ are not reflected by a similar amount of variation in $\Delta_{S}$, and these two quantities match each other best at sites lying within clusters of multiple patches with similar values of $\Delta_{T}$ are present. As for $\Delta_{L}$, we find that not only is its variation much smaller than either $\Delta_{T}$ or $\Delta_{S}$, but its variations are also much more long-ranged than the other two quantities: while $\Delta_{T}$ and $\Delta_{S}$ vary on the scale set by $l$, $\Delta_{L}$ on the other hand varies over longer length scales (several multiples of $l$). The sharp boundaries are largely invisible to $\Delta_{L}$, which only curves slightly in the presence of a boundary between small- and large-$\Delta_{S}$ regions. We also find that $\Delta_{L}$ tends to diverge from $\Delta_{T}$ or $\Delta_{S}$ most strongly when $\Delta_{T}$ is large; in this instance, it tends to remain near $\Delta_{L}\approx\overline{\Delta}$, and only increases mildly in the presence of this strong-pairing patch.
@@
In sum, what we have observed here is that for the inhomogeneous $d$-wave superconductor with $l\approx\xi$, the low-energy ($\omega<\Delta_{K}$) spectra are homogeneous (for the $l=8$ results here, the low-energy spectra are that of a homogeneous $d$-wave superconductor with $\Delta=0.16$), while the high-energy ($\omega>\Delta_{K}$) LDOS exhibits much of the full inhomogeneity of the underlying order parameter, but with uneven correspondences between $\Delta_{S}$ and $\Delta_{T}$. Notably, proximity effects are strong; this is most evident in the spectra with $\Delta_{T}\approx 0$, which show a finite $\Delta_{S}$. The critical energy $\Delta_{K}$, which separates homogeneous and inhomogeneous states, in turn appears to be largely set by the smallest spectral gaps in the system. A crucial difference between the $l=8$ regime studied here and the $l=1$ one in the previous section is that the quasiparticle response of the system to order-parameter inhomogeneity largely hews to the underlying order parameter at high energies such that $\Delta_{T}$ and $\Delta_{S}$ are positively correlated, but at low energies the response reflects little of the order-parameter variation (so that $\Delta_{L}$ is only weakly correlated with $\Delta_{T}$). Evidently, increasing the inhomogeneity length scale $l$ reduces the energy range in which the spectrum is homogeneous, so that for sufficiently large $l$, the resulting critical energy scale $\Delta_{K}$ separating homogeneous and inhomogeneous spectra is pushed downwards towards the Fermi energy.
@@ V Large Patches ($l>\xi$)
Figure 13: Plots of $\Delta_{T}$ (a), $\Delta_{S}$ (c), and $\Delta_{L}$ (e), and the absolute value of their respective Fourier transforms (b, d, and f) for patch size $l=32$. Only one disorder realization is presented. Shown here are quantities taken from the middlemost $256\times 256$ segment of the full $1024\times 256$ system.
Figure 14: Top left: Plots of the binned average local density of states for $l=32$. Each LDOS spectrum is binned according to its spectral gap (with 10 bins in all), and all spectra in a given bin are then averaged over. Top right: Plot of the distribution of the local density of states as a function of energy (heat map) and the average local density of states (dashed white line) for $l=32$. The spectra are binned according to the energy, and histograms for each energy are taken. The counts per energy bin are shown as a heat map. For both these plots, 4 disorder realizations and a total of 1,048,576 individual spectra are used in this plot. Bottom: Plots of the density of states for a clean bulk $d$-wave superconductor for various values of $\Delta$. To obtain these plots, we made use of a $1024\times 1024$ momentum-space grid, and the same band-structure and broadening parameters in the inhomogeneous cases are used here.
@@
In this regime, both $\Delta_{T}$ and $\Delta_{S}$ maps look mostly identical to one another; this can be seen in Fig. 13a and c. The difference is primarily in the more rounded shape of the patches in $\Delta_{S}$, and in the fact that when adjacent patches in $\Delta_{T}$ have similar (but not identical) values, they show up in $\Delta_{S}$ as a large, mostly uniform patch, with the gradations in $\Delta_{T}$ not visible in $\Delta_{S}$. Also, isolated small-$\Delta_{S}$ regions tend to be slightly larger in spatial extent than similarly isolated large-$\Delta_{S}$ regions, an observation that is also clearly seen in the $l\approx\xi$ case; this is a reflection of the larger local coherence length within these small-gap patches, which ensures that (through the proximity effect) the effect of the small-$\Delta_{T}$ patch “leaks out” into the surrounding large-$\Delta_{T}$ region. Conversely, a large-$\Delta_{T}$ patch surrounded by a small-$\Delta_{T}$ would feature the “leakage” of the small-gap background into the patch, making the large-$\Delta_{S}$ patch look smaller than its actual size in $\Delta_{T}$ is. This effect for $l>\xi$ however is much more muted compared to what is seen in the $l\approx\xi$ case. Finally, similar to the $l=1$ and $l=8$ cases discussed earlier, the $\Delta_{S}$ map always is nonzero, with a minimum $\Delta_{S}$ present even in regions where $\Delta_{T}$ is close to zero. Even with $l=32$, proximity effects are strong enough that a nonzero pair density, and consequently a finite gap, is present deep within these patches. This minimum gap however is smaller than that seen in the $l\approx\xi$ case.
@@
The map of the low-energy gap $\Delta_{L}$ with $l=32$ (Fig. 13e) now clearly reflects the inhomogeneous nature of the underlying order parameter. Small- and large-$\Delta_{T}$ regions match up neatly with small- and large-$\Delta_{L}$ regions, and this time the distribution of $\Delta_{T}$ is broad enough to capture the variations in $\Delta_{T}$ to reasonable accuracy. The agreement is clearly not perfect, though: the finer gradations seen in $\Delta_{T}$ are not captured by $\Delta_{L}$, and the distribution of values of $\Delta_{L}$ still appears to be narrower than that of $\Delta_{S}$. Nevertheless, the $\Delta_{L}$ plots demonstrate that for $l>\xi$, the low-energy states now display more inhomogeneity than those in the $l\approx\xi$ and $l<\xi$ cases, but still not as much as the high-energy states do. The overall similarity among the $\Delta_{T}$, $\Delta_{S}$, and $\Delta_{L}$ maps carries over to their respective Fourier transforms (Figs. 13b, d and f). In particular, the Fourier transforms of the $\Delta_{T}$ and $\Delta_{S}$ maps look almost identical, with both featuring a highly localized region of high intensity near zero wavevector and higher harmonics arising from the regular square shape of the patches in both real-space maps. The Fourier transform of the $\Delta_{L}$ map also features strong intensities only at small wavevectors. One can see that the small-momentum part of Fig. 13f is similar to that of Figs. 1f and 7f, but as higher-momenta scattering processes are suppressed for the $l=32$ case, the only feature in the Fourier transform left is a much smaller X-shaped feature near zero momentum.
@@
Similar to the $l=8$ case, at $\omega=0.16$ (Fig. 15c), there are three kinds of regions: large-LDOS patches which correspond to patches where the local spectral gap $\Delta_{S}\approx 0.16$; small-LDOS patches which correspond to large-$\Delta_{S}$ regions with $\Delta_{S}>0.16$; and regions hosting prominent modulations which correspond to small-$\Delta_{S}$ regions with $\Delta_{S}<0.16$. These striking modulations are (like those seen in the $l=8$ case) a result of quasiparticle scattering off of the walls of the order-parameter barrier formed by the large-$\Delta_{T}$ neighboring patches. These regions with modulations grow in size with increasing $\omega$. At $\omega=0.31$ (Fig. 15d), the only regions without any visible modulations are the large-LDOS patches corresponding to patches with large $\Delta_{S}$; these modulations are present everywhere else. A peculiarity of our square-patch disorder model is that when $l$ is large (such as in the present $l=32$ case), the boundaries of the small-$\Delta_{S}$ regions that host these modulations are straight and have right-angle corners, which result in mostly ordered modulations in the $(1,0)$ and $(1,0)$ directions. This is in contrast with that seen in the $l=8$ case earlier where, as a result of the small value of $l$, the regions that have these modulations are irregularly shaped and therefore the resulting scattering patterns are not as sharply defined as in the present case.
@@
To succinctly summarize all these observations together, we show once more the various two-dimensional histograms of the order-parameter measures in Fig. 17. The corresponding plot for the $\Delta_{T}$-$\Delta_{S}$ pair (Fig. 17a) shows that the two quantities are now very strongly correlated with one another, with only minimal fuzziness in the correspondences. It is worth noting that the system still exhibits a finite $\Delta_{S}$ everywhere, even in regions where $\Delta_{T}\approx 0$. Here, the smallest $\Delta_{S}\approx 0.05$, which is smaller than the smallest $\Delta_{S}$ for the $l\approx\xi$ case. This is consistent with the observation made earlier that the kink energy $\Delta_{K}$, which is the scale that separates homogeneous and inhomogeneous states and which looks to be tied to the smallest spectral gaps in the system, is pushed down in energy as the inhomogeneity length scale $l$ is increased. The histogram for the $\Delta_{T}$-$\Delta_{L}$ pair (Fig. 17b) now shows a much more pronounced positive correlation between these two quantities compared with the two cases considered before, with the agreement best at small $\Delta_{T}$ and worst at large $\Delta_{T}$. Finally, the $\Delta_{S}$-$\Delta_{L}$ pair (Fig. 17c) also appears to be more correlated with one another than in the previous cases. As with the $l\approx\xi$ case, here we find a similar tendency for the joint distribution to be mostly localized within one small region for small $\Delta_{S}$ and for it to extend out for increasing $\Delta_{S}$.
@@ VI Low-Energy Homogeneity and Spectral Kinks
Figure 19: Plots of the binned average local density of states for $l=1$ (a), $l=8$ (b), and $l=32$ (c), with finer bins used than in the previous plots. Each LDOS spectrum is binned according to its spectral gap (with 41 bins in all), and all spectra in a given bin are then averaged over. For these plots, 4 disorder realizations and a total of 1,048,576 individual spectra are used. The width of each line plot is proportional to the number of spectra within the bin. Only lineshapes corresponding to bins with total number of spectra greater than 1024 are shown here.
@@
Perhaps the most notable behavior seen in STS experiments on the cuprates is the sharp dichotomy between the low-energy homogeneous states and the high-energy inhomogeneous states that are characteristic of the $l\gtrapprox\xi$ case, which is accompanied by the presence of kinks appearing in large-$\Delta_{S}$ spectra. We had noticed in Figs. 8a and 14a that these kinks tend to appear at a characteristic kink energy $\Delta_{K}$ which is roughly set by the smallest spectral gaps of the system. In this section, we define $\Delta_{K}$ appropriately and study it as a function of $l$, contrast its behavior with the other energy scales that can be obtained from the LDOS spectra ($\Delta_{S}$ and $\Delta_{L}$), and put on quantitative footing the informal observations made in the previous section.
We first define $\Delta_{K}$ in quantitative terms. One can observe in Figs. 10a and 16a that a purely local definition of $\Delta_{K}$ as the energy where a kink is present is difficult to implement in practice in our numerics: at low energies, numerous wiggles in the LDOS as a function of energy can be seen, and it is difficult to ascertain if these are the true kinks that separate the homogeneous states from the inhomogeneous ones, or if these arise from finite-size effects. This difficulty of isolating the kink in the spectrum is otherwise not present in our operational definitions of $\Delta_{S}$ (which involves simply finding the peak of the positive-energy LDOS) and $\Delta_{L}$ (which involves taking the slope of the LDOS near $\omega=0$).
To get around this difficulty, we work instead with the bin-averaged spectra, as visualized in Fig 19. Here, we bin the individual spectra by its local spectral gap. Unlike Figs. 2a,  8a, and 14a, however, here we use much smaller bin widths that allow us to capture the averaged spectrum for each value of $\omega$. Because of the very large number of individual spectra, even a finer binning procedure allows for enough spectra per bin to ensure that an average over a sufficiently large number of spectra is performed for each individual bin. The averaging over a large number of spectra allows us to smoothen out any finite-size effects, and we can thus define $\Delta_{K}$, as well as $\Delta_{S}$ and $\Delta_{L}$, for each of the bin-averaged spectra. $\Delta_{K}$ is defined as the lowest energy at which the derivative with respect to $\omega$ of the positive-energy bin-averaged LDOS exhibits a peak. Note that this definition implies that $\Delta_{K}\approx\Delta_{S}$ in the absence of a kink. Right away, we can expect the following behaviors to hold at these two extreme limits: a) in the extreme pointlike limit $l\ll\xi$ where the states are highly homogeneous, all three energy scales coincide, i.e., $\Delta_{S}\approx\Delta_{L}\approx\Delta_{K}\approx\text{constant}$; and b) in the extreme bulk limit $l\to\infty$, $\Delta_{S}\approx\Delta_{T}$ while $\Delta_{K}\to 0$.
@@
These results suggest that in the physically relevant $l\approx\xi$ regime, the notion of an average spectral gap—defined either by taking both the system- and configuration-average of $\Delta_{S}$ or by binning the LDOS and obtaining the bin-averaged $\Delta_{S}$, as we have done here—leads to a considerable overestimate of the true average underlying order parameter $\overline{\Delta}$. This is partly due to the finite $\Delta_{S}$ even in regions where $\Delta_{T}\approx 0$: these finite $\Delta_{S}$ values push up the average, leading to an overestimate of $\overline{\Delta}$. As we had noted earlier, these particular $\Delta_{S}$ values (i.e., the smallest gaps) are correlated with the value of $\Delta_{K}$; therefore, as $l$ increases, $\Delta_{K}$, and therefore these particular $\Delta_{S}$’s, decrease toward zero, and the average of $\Delta_{S}$ regresses to $\overline{\Delta}$ for sufficiently large $l$. It is telling that in Fig. 21, even patches as large as $l=64$ for the $\overline{\Delta}=0.32$ case still feature a large enough discrepancy between the averages of $\Delta_{S}$ and $\Delta_{L}$ to indicate that the system in this case is still far from the extreme bulk limit. On the other hand, taking the average of $\Delta_{L}$ appears to give a more accurate estimate of $\overline{\Delta}$, owing to the robustness of low-energy homogeneity when $l\approx\xi$.